# Large Language Models Hallucination: A Comprehensive Survey

### 09 Oct 2025 v2

@@ 6.5 Mitigation Challenges
[124] 2024 Instruction generation (Legal) Private Acc., Fluency
[125] 2023 QA citation Private Accu.
@@
[134] 2023 QA DROP, GSM-IC, EntailmentBank Acc.
[136] 2023 QA MS MARCO Acc., human
[140] 2024 QA NQ, TriviaQA, SQuAD-1, BioASQ Coverage rate
[142] 2024 QA TruthfulQA, StrategyQA, NQ, TriviaQA GPT-judge, BLEU, Rouge, Acc.
[145] 2024 QA OBQA,CSQA, WQSP, CWQ Acc., Hits@1
@@
[150] 2023 QA WebQuestionsSP, Mintaka MRR, Top-K Acc.
[152] 2023 QA HotpotQA, 2WikiMultihopQA, MuSiQue, IIRC Recall, F1
[155] 2024 QA Simple Question, Mintaka, HotpotQA EM, F1
[132] 2023 QA, Summarization Private Rouge, Human
[133] 2023 QA, Dialogue WikiQA, DSTC7/11
[135] 2023 QA, Dialogue NQ, SQA, QReCC Human, Attrauto, PresLev
[137] 2024 QA, Summarization, Data-to-Text RAGTruth Precision, Recall, F1, Human
[157] 2024 Dialogue DoQA, QuAC F1, Acc., Faithfulness
[158] 2025 Code generation HumanEval, MBPP, MBCPP Pass@k
[171] 2024 QA HotpotQA, 2WikiMultiHopQA F1, human
[172] 2024 QA KNOWREASON (Wikidata5m), CommonsenseQA, ARC, BBH EM
[173] 2023 QA FEVER, HotpotQA, FeTaQA, MedMCQA, MMLU (Phys/Bio) Acc., EM
[174] 2024 QA NQ, TriviaQA, WebQ, RealTimeQA F1, EM
[205] 2024 QA, Text generation Wikidata, Wiki-Category List, MultiSpanQA, Biographies F1, Precision, Recall, Factscore
[170] 2023 QA, Summarization HaluSum2130, HaluQA4170, FactCC503, SummEval, QAGS-CNNDM/XSUM F1
[161] 2024 Multiple tasks Multiple datasets Acc.
[177] 2023 Machine translation FLORES-101, WMT COMET
[178] 2024 Machine translation HLMT, X-branch, High-res SpBLEU, ChrF2
[176] 2024 Summarization CNN-DM, XSUM ROUGE-L, BERT-Precision, FactKB
[181] 2025 Summarization CNNDM, XSUM ROUGE-L, FC, FR
[190] 2024 Summarization XSum, CNN/DM QuestEval, FactCC, Rouge, Human
[195] 2023 Summarization XLSum mFact, Rouge, Human
[175] 2023 QA TruthfulQA, FACTOR, StrategyQA, GSM8K, Vicuna QA Acc., LLM-as-a-judge
[182] 2024 QA SVMAP, CQA Acc.
[183] 2023 QA Private HALOCHECK, Human
[184] 2024 QA HaluEval, TruthfulQA, FACTOR Acc.
[194] 2025 QA CRAFT, Alpaca Acc., Consistency
[191] 2023 Dialogue WoW F1, Rouge, BLEU, Meteor, KF1, EF1 Acc., Human
[188] 2024 Multiple tasks Multiple datasets Acc., F1, EM
@@ 7 Benchmark Datasets
Ref Name Year Lang. Domain Source Size Detection Mitigation GT Source
@@ 9.1 Hallucination Detection
Lack of Explainability. Some detection methods do not provide explanations alongside their predictions. HuDex [231] integrated explainability with hallucination detection. However, the generated explanations must be faithful and concise, which requires bridging the gap between model-internal metrics and human reasoning. Future research should concentrate on producing comprehensible explanations for the rationale behind categorizing the response as hallucinated. This would enhance user confidence and facilitate model debugging.

### 18 Mar 2026 v3

@@ 7.5 Mitigation Challenges
The observed limitations across mitigation strategies indicate that no single approach can address all hallucination failure modes, underscoring the need for domain-aware, task-adaptive pipelines that combine complementary techniques. Retrieval-based approaches tend to be more beneficial in knowledge-intensive domains, such as medicine, law, and finance. However, retrieval-based mitigation is less reliable in low-resource languages or rapidly evolving domains, where trustworthy sources are scarce or inconsistent. In contrast, reasoning-based strategies are more helpful for multi-step reasoning tasks, such as math logical QA, but may not be useful in domains where hallucinations arise from missing or incorrect world knowledge rather than from flawed reasoning chains. Prompt-based and decoding-based mitigation techniques often show gains in structured or well-defined tasks, such as summarization or information extraction, where instructions can more tightly control faithfulness to input. However, their effectiveness declines in open-ended generation settings, where hallucinations arise from broad knowledge gaps rather than instruction-following failures. Similarly, model-centric approaches such as fine-tuning, contrastive learning, and distillation tend to perform well in domain-specific deployments with curated training data, but they struggle to generalize to unseen domains and may even amplify domain-specific biases or misconceptions. This highlights the need for domain-aware, task-adaptive mitigation pipelines that combine retrieval, reasoning control, and model-level adaptation tailored to the deployment setting.
@@
[148] 2024 Legal instruction gen Private Accuracy, Fluency
[149] 2023 QA citation Private Accuracy
@@
[158] 2023 QA DROP, GSM-IC, EntailmentBank Accuracy
[160] 2023 QA MS MARCO Accuracy, Human
[164] 2024 QA NQ, TriviaQA, SQuAD, BioASQ Coverage rate
[166] 2024 QA TruthfulQA, StrategyQA, NQ, TriviaQA GPT-judge, BLEU, Rouge
[169] 2024 QA OBQA, CSQA, WQSP, CWQ Accuracy, Hits@1
@@
[174] 2023 QA WebQuestionsSP, Mintaka MRR, Top-K Accuracy
[177] 2023 QA HotpotQA, 2WikiMultihopQA Recall, F1
[181] 2024 QA Mintaka, HotpotQA EM, F1
[176] 2025 QA Follow-up QA Accuracy, EM
[179] 2025 QA CWQ, WebQSP, GrailQA EM
[159] 2023 QA, Dialogue NQ, SQA, QReCC Human, Attrauto
[161] 2024 QA, Summarization RAGTruth Precision, Recall, F1
[183] 2024 Dialogue DoQA, QuAC F1, Accuracy
[199] 2024 QA HotpotQA F1, Human
[200] 2024 QA CommonsenseQA, ARC EM
[201] 2023 QA FEVER, HotpotQA Accuracy, EM
[202] 2024 QA NQ, TriviaQA F1, EM
[196] 2025 QA TruthfulQA, HaluEval Accuracy
[205] 2023 Machine translation FLORES-101 COMET
[206] 2024 Machine translation HLMT SpBLEU, ChrF2
[204] 2024 Summarization CNN-DM, XSUM ROUGE-L, FactKB
[210] 2025 Summarization CNNDM, XSUM ROUGE-L
[221] 2024 Summarization XSum, CNN/DM QuestEval, FactCC
[226] 2023 Summarization XLSum mFact, Rouge
[203] 2023 QA TruthfulQA Accuracy
[212] 2023 QA Private HALOCHECK
[213] 2024 QA HaluEval, TruthfulQA Accuracy
[207] 2025 QA TruthfulQA, TriviaQA Accuracy
[217] 2025 QA TruthfulQA, TriviaQA Accuracy
[222] 2023 Dialogue WoW F1, Rouge, BLEU
[219] 2024 Multiple tasks Multiple datasets Accuracy
@@ 8 Benchmark Datasets
Tables II and V show that hallucination detection and mitigation remain highly task- and benchmark-dependent. No single detection or mitigation method consistently outperforms others across all datasets. The performance of detection and mitigation techniques fluctuates under domain and task shifts. Furthermore, the diversity of benchmarks and evaluation setups complicates direct comparisons, which underscores the need for standardized evaluation protocols and hybrid approaches tailored to specific task characteristics.
While most existing benchmarks focus on evaluating the performance of hallucination detection and mitigation methods rather than measuring how frequently hallucinations occur in LLMs, recent community efforts provide useful empirical data on hallucination rates across systems. One such resource is the Hallucination Leaderboard [235], which aggregates model performance on grounded summarization tasks and reports hallucination rates computed by the Hughes Hallucination Evaluation Model (HHEM) [235]. Some state-of-the-art models achieve hallucination rates below a few percentage points on this task. This indicates that in narrow, constrained settings, such as summarization, hallucinations are relatively infrequent compared with unconstrained generation. However, substantial variation remains across model families and prompt behaviors, which demonstrates that hallucination is neither uniform nor negligible even under strict grounding conditions. While such task-specific benchmarks do not yet provide a unified, cross-task hallucination frequency metric, the emerging leaderboard results represent one of the first systematic efforts to quantify hallucination prevalence and offer the community a common reference point for comparing models on factual consistency.
@@
Ref Name Year Lang. Detect Mitigate Domain Source Size GT Source
@@
TABLE V: Benchmark results for hallucination mitigation techniques using the most utilized datasets.
Ref TriviaQA NQ HotpotQA XSum TruthfulQA CNN/DM
EM F1 EM F1 RL EM F1 RL RL Acc RL
[152] 17.6 26.9
[166] 69.0 78.5 40.0 57.3 31.2
[177] 49.3 60.7
[181] 34.0 47.2
[159] 57.0
[199] 90.3
[201] 35.4
[204] 20.0 27.1
[210] 7.4 20.0 27.8
[221] 38.2
[207] 80.8 47.4
[196] 45.1
[217] 48.4
[207] 31.0 50.2
[202] 76.3 82.6 57.5 48.9 50.2
@@ 10.3 Benchmarks and Evaluation Metrics
Bias of LLM-as-Judge. Recent evidence shows that LLM-as-a-judge can be influenced by visually appealing formatting, which leads to large preference shifts even when the underlying content quality is unchanged [268]. This is particularly concerning for open-ended evaluation where judges may be convinced by perceived credibility rather than factual correctness [268]. In addition, LLM-as-a-judge exhibits positional effects. The same judge may prefer different candidates when the order of candidates is swapped, and this behavior is not explained by random variation [269]. LLM-as-a-judge also suffers from cognitive-related biases. Therefore, it is recommended to avoid using the same model to both generate and judge outputs [270]. Future research should investigate using jury-style aggregation across multiple diverse judges to reduce self-enhancement and idiosyncratic preferences [270, 271]. It is also recommended to test the judges using format perturbations, fake-reference perturbations, and adversarial phrasing tests as part of the evaluation protocol.